# Communication-Efficient and Privacy-Adaptable Mechanism for Federated Learning

### v2

@@ I Introduction and Motivation
Federated learning (FL) [1] enables the training of machine learning models using vast amounts of data held privately by edge devices by exploiting their computational capabilities [2], thanks to its decentralized design [3]. However, FL presents challenges [4], such as ensuring efficient communication during training phases and preserving the privacy of the sensitive data at edge devices. These challenges are commonly addressed separately through methods that introduce some distortion to updated models, like compression methods [5, 6, 7, 8, 9] and differential privacy (DP) mechanisms [10, 11].
@@
To address both challenges simultaneously, a common approach involves initially applying a DP mechanism (such as Laplace [10] or Gaussian [11]) followed by quantization [12, 13]. However, this approach has two drawbacks. Firstly, the overall error (comprising privacy-preserving and quantization errors) will no longer be exact, leading to a degradation in the accuracy of model updates and the final trained model. Secondly, the quantization noise does not contribute to enhancing privacy against the honest but inquisitive clients within the FL framework. In essence, the DP-then-quantize approach is suboptimal.
@@
An alternative approach relies on randomized quantization, such as subtractive dithering [14, 15, 16, 17, 18]. Previous works [19, 20, 21, 22, 23] have employed this method to develop a joint privacy and quantization mechanism for FL. However, these approaches are limited to scalar quantization, which is less effective than vector quantization in terms of compression ratio. Moreover, Lang et al. [24] investigated the use of subtractive dithering to generate privacy-preserving noises such as multivariate $t$-distributed and Laplace noises for local DP. However, their proposed mechanism did not encompass the Gaussian mechanism. [^1] The mechanism proposed in [24] adds a carefully designed, low-power privacy-preserving noise to the model update before the dithered quantization step. However, the authors do not specify how to design the requisite infinitesimal noise to generate the Gaussian noise for local DP, a task that is highly non-trivial. In contrast, our proposed CEPAM specifies how to generate Gaussian noise for central DP.
@@
Channel simulation [25, 26, 27, 28, 29, 30, 31, 32] is technique that can achieve the same goal. Recognizing that the Gaussian distribution can be represented as a mixture of uniform distributions [33, 34], Agustsson et al. [35] combined subtractive dithering with layered construction [36] to simulate 1D Gaussian channel. This approach was further developed in [37] to simulate any 1D unimodal additive noise distribution using randomized scaling and offsets. Recently, a new construction in [38] named RSUQ, which is a randomized vector quantization, was introduced to exactly simulate any multivariate nonuniform continuous additive noise distribution with finite communication by extending 1D methods to multivariate distributions. However, [38] did not consider FL setup.
@@
In this work, we propose a joint mechanism, CEPAM, aimed at enhancing privacy while compressing model updates in FL. Our proposed CEPAM leverages RSUQ [38], building upon its capability to convert quantization distortion into an additive noise term with adjustable variance, independent of the quantized data. Within CEPAM, both clients and the server can tailor the privacy protection mechanism, offering the choice between the exact Gaussian mechanism (CEPAM-Gaussian) or Laplace mechanism (CEPAM-Laplace), based on the required accuracy level and privacy protection. Additionally, the inherent nature of RSUQ as a vector quantizer gives CEPAM the advantage in reducing the compression ratio compared to its scalar counterparts [21, 22, 35, 36, 37].
We theoretically analyze the privacy guarantee of CEPAM and investigate the trade-offs among user privacy and accuracy of CEPAM through experimental evaluations. Additionally, we analyze the privacy of CEPAM-Gaussian and CEPAM-Laplace using techniques from privacy amplification [39, 40]. We validate our theoretical results by comparing CEPAM-Gaussian against other baselines using the MNIST dataset with MLP and CNN architectures. The experimental results demonstrate that for MLP and CNN models, CEPAM-Gaussian achieves improvements of 0.6-2.0% and 0.4-1.0% in accuracy, respectively, compared to other baselines.
@@ II Related Works
The channel simulation problem [25, 41, 42] is closely related to Shannon’s lossy compression problem. In this setup [25], an encoder compresses an input source $X$ into a message $M$, which a decoder then uses to generate an output $M$. The encoder and decoder may share common randomness. The objective is to ensure that $Y$ follows a prescribed conditional distribution given $X$. Channel simulation was initially studied in the asymptotic setting [25, 41, 42], where $X$ is a sequence, and has also been investigated in the one-shot setting [26, 43, 27, 44], where $X$ is a scalar. RSUQ [38] can be regarded as a one-shot channel simulation scheme for simulating an additive noise channel with an arbitrary continuous noise distribution. For a comprehensive treatment of the subject, we direct the interested reader to the monograph in [32].
@@
Federated learning (FL) was initially introduced by McMahan et al. [1] as a collaborative learning approach that allows model training without the need to gather user-specific data. Research in FL focuses mainly on improving communication efficiency [1, 5, 8, 45, 46, 47, 9, 48, 49], ensuring data privacy [50, 51], or an approach that addresses both aspects [21, 22, 31]. This work is focused on an approach that addresses the issues of communication efficiency and data privacy simultaneously.
@@
The communication cost in federated learning can be mitigated through two primary methods: reducing the number of communication rounds or compressing model updates (e.g., parameter vectors or stochastic gradients) before transmission. While the original FedAvg algorithm [1] operates heuristically under heterogeneous data (i.e., non-iid data) and device inactivity (i.e., stragglers), Local SGD explicitly reduces communication frequency by performing multiple local update steps before synchronization. However, the convergence guarantees for Local SGD, which were first established by Stich [47] for strongly-convex and smooth objectives, relied on the idealized assumptions of iid data and no stragglers. Li et al. [49] later extended this analysis by providing theoretical guarantees for Local SGD in settings with heterogeneous data and inactive devices. Subsequently, Shlezinger et al. [9] integrated compression into this framework with UVeQFed, which incorporates quantization for heterogeneous settings.
@@
Another major method to reducing communication costs is the compression of stochastic gradients prior to transmission [8, 46, 9]. This includes methods like quantization, which constrains the number of bits used per value. A key technique is unbiased quantization [8], which ensures the compressed gradients remain unbiased estimators of their true values. In contrast, sparsification techniques [45, 46] improve communication efficiency by reducing the number of non-zero entries in the stochastic gradients. Building on these advances, we propose a novel Local SGD variant that unifies quantization with a privacy-preserving mechanism based on RSUQ, simultaneously addressing communication efficiency and data confidentiality within a single framework.
@@
A recent trend involves developing approaches that jointly address communication efficiency and privacy [19, 29, 21, 22, 24, 23, 31, 52]. For central DP, [21, 22] used subtractive dithering to achieve this for scalar entries. CEPAM generalizes the scalar counterparts in [21, 22], as it uses RSUQ, which is a vector quantizer. For local DP, [24] proposed injecting low-power noise before dithered quantization step, proving guarantees for Laplace distribution and multidimensional $t$-distribution. However, the authors do not specify how to generate the requisite infinitesimal noise to produce Gaussian noise. In contrast, our CEPAM directly specifies the generation of this Gaussian noise for central DP. In parallel, techniques from channel simulation have inspired other DP compressors [29, 31].
@@ III-A Federated Learning (FL)
In this work, we consider the FL framework (or federated optimization) proposed in [1]. More explicitly, $K$ clients (or devices), each attached with a local dataset $\mathcal{D}^{(k)}$ where $k\in\{1,2,...,K\}=:\mathcal{K}$, cooperate together to train a shared global model $\mathbf{W}$ with $m$ parameters through a central server. The goal is to minimize the objective function $F:\mathbb{R}^{m}\to\mathbb{R}$:
@@
$F_{k}(\mathbf{W}){\color[rgb]{0,0,0}\equiv F_{k}(\mathbf{W},\mathcal{D}^{(k)})}:=\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}\ell(\mathbf{W};\xi_{k,j}),$ (2)
@@
Let $T$ denote the total number of iterations in FL and let $\mathcal{T}_{T}:=\{0,\tau,2\tau,\ldots,T\}$ denote the set of integer multiples of some positive integer $\tau\in\mathbb{Z}^{+}$ where $T\equiv 0\pmod{\tau}$, called the set of synchronization indices. We describe one FL round of the conventional FedAvg [1] for solving optimization problem (1) as follows. Let $\mathbf{W}_{t}$ denote global parameter vector available at the server at the time instance $t\in\mathcal{T}_{T}$. At the beginning of each FL round, the server broadcasts $\mathbf{W}_{t}$ to all the clients. Then, each client $k$ sets $\mathbf{W}_{t}^{k}=\mathbf{W}_{t}$ and computes the $\tau\;(\geq 1)$ local parameter vectors by SGD: [^2] When $T$ is fixed, the larger $\tau$ is, the fewer the communication rounds.
$\mathbf{W}_{t+t^{\prime}}^{k}\leftarrow\mathbf{W}_{t{\color[rgb]{0,0,0}+t^{\prime}-1}}^{k}-\eta_{t+t^{\prime}}\nabla F_{k}^{j_{t{\color[rgb]{0,0,0}+t^{\prime}-1}}^{k}}(\mathbf{W}_{t{\color[rgb]{0,0,0}+t^{\prime}-1}}^{k}),\;\;t^{\prime}=1,\ldots,\tau,$ (3)
where $\eta_{t+t^{\prime}}$ is the learning rate, $\nabla F_{k}^{j}(\mathbf{W}):=\nabla F_{k}(\mathbf{W};\xi_{k,j})$ is the gradient computed at a single sample of index $j$, and $j_{t}^{k}$ is the sample index chosen uniformly from the local data $\mathcal{D}^{(k)}$ of client $k$ at time tt. [^3] In this work, our focus is on analyzing the computation of a single stochastic gradient at each client during every time instance. The FL convergence rates can potentially be enhanced by incorporating mini-batching technique [47], leaving the detailed analysis for future work. Finally, by assuming that all clients participate in each FL round for simplicity, the server aggregates the $K$ local model updates $\{\mathbf{X}_{t+\tau}^{k}=\mathbf{W}_{t+\tau}^{k}-\mathbf{W}_{t}\}_{k\in\mathcal{K}}$ and computes the new global parameter vector:
@@
In [9], Shlezinger et al. introduced a framework for Quantized Federated Learning. Given the limited bandwidth of uplink channel in FL framework, each $k$-th client is required to communicate a quantized version [53] of the model update $\mathbf{X}_{t+\tau}^{k}$ using a finite number of bits to the server. Specifically, the $k$-th model update $\mathbf{X}_{t+\tau}^{k}$ is encoded into a binary codeword $u_{t+\tau}^{k}$ of length $R_{k}$ bits through an encoding function $\mathrm{Enc}_{t+\tau}^{k}:\mathbb{R}^{m}\to\{0,1,\ldots,2^{R_{k}}-1\}:=\mathcal{U}_{k}$. Since the outputs $\mathrm{Enc}_{t+\tau}^{k}(\mathbf{X}_{t+\tau}^{k})=u_{t+\tau}^{k}\in\mathcal{U}_{k}$ of an encoder may have unequal probabilities, entropy coding [54, 55, 56] can be utilized to further reduce redundancy when transmitting through the binary lossless channel. In the FL literature, the uplink channel is typically modeled as a bit-constrained link [9], and the transmission is assumed to be lossless. Upon receiving the set of codewords $\{u_{t+\tau}^{k}\}_{k\in\mathcal{K}}$ from clients, the server uses the joint decoding function $\mathrm{Dec}_{t+\tau}:\mathcal{U}_{1}\times\ldots\times\mathcal{U}_{K}\to\mathbf{R}^{m}$ to reconstruct $\hat{\mathbf{X}}_{t+\tau}\in\mathbb{R}^{m}$, an estimate of the weighted average $\sum_{k\in\mathcal{K}}p_{k}\mathbf{X}_{t+\tau}^{k}$.

### v3

@@ I Introduction and Motivation
Federated learning (FL) [McMahan2016FL] enables the training of machine learning models using vast amounts of data held privately by edge devices by exploiting their computational capabilities [Chen2019DLEdgeCompute], thanks to its decentralized design [Jeffrey2012distributed]. However, FL presents challenges [Tian2019FLChallenges], such as ensuring efficient communication during training phases and preserving the privacy of the sensitive data at edge devices. These challenges are commonly addressed separately through methods that introduce some distortion to updated models, like compression methods [Konecn2016NIPSFederatedLS, lin2020DGCreducing, Corentin2017DLadapCompr, alistarh2017QSGD, shlezinger2020uveqfed] and differential privacy (DP) mechanisms [Dwork06DP, Dwork14Book].
@@
To address both challenges simultaneously, a common approach involves initially applying a DP mechanism (such as Laplace [Dwork06DP] or Gaussian [Dwork14Book]) followed by quantization [girgis2021shuffled, agarwal2021skellam]. However, this approach has two drawbacks. Firstly, the overall error (comprising privacy-preserving and quantization errors) will no longer be exact, leading to a degradation in the accuracy of model updates and the final trained model. Secondly, the quantization noise does not contribute to enhancing privacy against the honest but inquisitive clients within the FL framework. In essence, the DP-then-quantize approach is suboptimal.
@@
An alternative approach relies on randomized quantization, such as subtractive dithering [roberts1962picture, schuchman1964dither, Limb1969visual, Jayant1972speech, sripad1977necessary]. Previous works [Kim2021FederatedLW, shahmiri2024communication, hasircioglu2023communication, yan2023layered, hegazy2024compression] have employed this method to develop a joint privacy and quantization mechanism for FL. However, these approaches are limited to scalar quantization, which is less effective than vector quantization in terms of compression ratio. Moreover, Lang et al. [lang2023joint] investigated the use of subtractive dithering to generate privacy-preserving noises such as multivariate $t$-distributed and Laplace noises for local DP. However, their proposed mechanism did not encompass the Gaussian mechanism. [^1] The mechanism proposed in [lang2023joint] adds a carefully designed, low-power privacy-preserving noise to the model update before the dithered quantization step. However, the authors do not specify how to design the requisite infinitesimal noise to generate the Gaussian noise for local DP, a task that is highly non-trivial. In contrast, our proposed CEPAM specifies how to generate Gaussian noise for central DP.
@@
Channel simulation [bennett2002entanglement, harsha2010communication, sfrl_trans, havasi2019minimal, shah2022optimal, flamich2023greedy, liu2024universal, Li2024ChannelSim] is technique that can achieve the same goal. Recognizing that the Gaussian distribution can be represented as a mixture of uniform distributions [WalkerUniformP1999, CHOY2003exppower], Agustsson et al. [agustsson2020universally] combined subtractive dithering with layered construction [wilson2000layered] to simulate 1D Gaussian channel. This approach was further developed in [hegazy2022randomized] to simulate any 1D unimodal additive noise distribution using randomized scaling and offsets. Recently, a new construction in [Ling2024arxivRSUQ] named RSUQ, which is a randomized vector quantization, was introduced to exactly simulate any multivariate nonuniform continuous additive noise distribution with finite communication by extending 1D methods to multivariate distributions. However, [Ling2024arxivRSUQ] did not consider FL setup.
@@
In this work, we propose a joint mechanism, CEPAM, aimed at enhancing privacy while compressing model updates in FL. Our proposed CEPAM leverages RSUQ [Ling2024arxivRSUQ], building upon its capability to convert quantization distortion into an additive noise term with adjustable variance, independent of the quantized data. Within CEPAM, both clients and the server can tailor the privacy protection mechanism, offering the choice between the exact Gaussian mechanism (CEPAM-Gaussian) or Laplace mechanism (CEPAM-Laplace), based on the required accuracy level and privacy protection. Additionally, the inherent nature of RSUQ as a vector quantizer gives CEPAM the advantage in reducing the compression ratio compared to its scalar counterparts [hasircioglu2023communication, yan2023layered, agustsson2020universally, wilson2000layered, hegazy2022randomized]. We theoretically analyze the privacy guarantee of CEPAM and investigate the trade-offs among user privacy and accuracy of CEPAM through experimental evaluations. Additionally, we analyze the privacy of CEPAM-Gaussian and CEPAM-Laplace using techniques from privacy amplification [Balle2018PrivAmpl, Feldman2018PrivAmpl]. We validate our theoretical results by comparing CEPAM-Gaussian against other baselines using the MNIST dataset with MLP and CNN architectures. The experimental results demonstrate that for MLP and CNN models, CEPAM-Gaussian achieves improvements of 0.6-2.0% and 0.4-1.0% in accuracy, respectively, compared to other baselines.
@@ II Related Works
The channel simulation problem [bennett2002entanglement, winter2002compression, cuff2013distributed] is closely related to Shannon’s lossy compression problem. In this setup [bennett2002entanglement], an encoder compresses an input source $X$ into a message $M$, which a decoder then uses to generate an output $M$. The encoder and decoder may share common randomness. The objective is to ensure that $Y$ follows a prescribed conditional distribution given $X$. Channel simulation was initially studied in the asymptotic setting [bennett2002entanglement, winter2002compression, cuff2013distributed], where $X$ is a sequence, and has also been investigated in the one-shot setting [harsha2010communication, braverman2014public, sfrl_trans, li2021unified], where $X$ is a scalar. RSUQ [Ling2024arxivRSUQ] can be regarded as a one-shot channel simulation scheme for simulating an additive noise channel with an arbitrary continuous noise distribution. For a comprehensive treatment of the subject, we direct the interested reader to the monograph in [Li2024ChannelSim].
@@
Federated learning (FL) was initially introduced by McMahan et al. [McMahan2016FL] as a collaborative learning approach that allows model training without the need to gather user-specific data. Research in FL focuses mainly on improving communication efficiency [McMahan2016FL, Konecn2016NIPSFederatedLS, alistarh2017QSGD, wangni2017gradSpar, stich2018sparsifiedSGD, stich2018local, shlezinger2020uveqfed, li2020FedProx, Li2020On], ensuring data privacy [Bonawitz2016SecureAggre, geyer2018DPFL], or an approach that addresses both aspects [hasircioglu2023communication, yan2023layered, liu2024universal]. This work is focused on an approach that addresses the issues of communication efficiency and data privacy simultaneously.
@@
The communication cost in federated learning can be mitigated through two primary methods: reducing the number of communication rounds or compressing model updates (e.g., parameter vectors or stochastic gradients) before transmission. While the original FedAvg algorithm [McMahan2016FL] operates heuristically under heterogeneous data (i.e., non-iid data) and device inactivity (i.e., stragglers), Local SGD explicitly reduces communication frequency by performing multiple local update steps before synchronization. However, the convergence guarantees for Local SGD, which were first established by Stich [stich2018local] for strongly-convex and smooth objectives, relied on the idealized assumptions of iid data and no stragglers. Li et al. [Li2020On] later extended this analysis by providing theoretical guarantees for Local SGD in settings with heterogeneous data and inactive devices. Subsequently, Shlezinger et al. [shlezinger2020uveqfed] integrated compression into this framework with UVeQFed, which incorporates quantization for heterogeneous settings.
@@
Another major method to reduce communication costs is the compression of stochastic gradients prior to transmission [alistarh2017QSGD, stich2018sparsifiedSGD, shlezinger2020uveqfed]. This includes methods like quantization, which constrains the number of bits used per value. A key technique is unbiased quantization [alistarh2017QSGD], which ensures the compressed gradients remain unbiased estimators of their true values. In contrast, sparsification techniques [wangni2017gradSpar, stich2018sparsifiedSGD] improve communication efficiency by reducing the number of non-zero entries in the stochastic gradients. Building on these advances, we propose a novel Local SGD variant that unifies quantization with a privacy-preserving mechanism based on RSUQ, simultaneously addressing communication efficiency and data confidentiality within a single framework.
@@
A recent trend involves developing approaches that jointly address communication efficiency and privacy [Kim2021FederatedLW, shah2022optimal, hasircioglu2023communication, yan2023layered, lang2023joint, hegazy2024compression, liu2024universal, lang2025olala]. For central DP, [hasircioglu2023communication, yan2023layered] used subtractive dithering to achieve this for scalar entries. CEPAM generalizes the scalar counterparts in [hasircioglu2023communication, yan2023layered], as it uses RSUQ, which is a vector quantizer. For local DP, [lang2023joint] proposed injecting low-power noise before the dithered quantization step, proving guarantees for Laplace distribution and multidimensional $t$-distribution. However, the authors do not specify how to generate the requisite infinitesimal noise to produce Gaussian noise. In contrast, our CEPAM directly specifies the generation of this Gaussian noise for central DP. In parallel, techniques from channel simulation have inspired other DP compressors [shah2022optimal, liu2024universal].
@@ III-A Federated Learning (FL)
In this work, we consider the FL framework (or federated optimization) proposed in [McMahan2016FL]. More explicitly, $K$ clients (or devices), each attached with a local dataset $\mathcal{D}^{(k)}$ where $k\in\{1,2,...,K\}=:\mathcal{K}$, cooperate together to train a shared global model $\mathbf{W}$ with $m$ parameters through a central server. The goal is to minimize the objective function $F:\mathbb{R}^{m}\to\mathbb{R}$:
@@
$F_{k}(\mathbf{W})\equiv F_{k}(\mathbf{W},\mathcal{D}^{(k)}):=\frac{1}{n_{k}}\sum_{j=1}^{n_{k}}\ell(\mathbf{W};\xi_{k,j}),$ (2)
@@
Let $T$ denote the total number of iterations in FL and let $\mathcal{T}_{T}:=\{0,\tau,2\tau,\ldots,T\}$ denote the set of integer multiples of some positive integer $\tau\in\mathbb{Z}^{+}$ where $T\equiv 0\pmod{\tau}$, called the set of synchronization indices. We describe one FL round of the conventional FedAvg [McMahan2016FL] for solving optimization problem (1) as follows. Let $\mathbf{W}_{t}$ denote global parameter vector available at the server at the time instance $t\in\mathcal{T}_{T}$. At the beginning of each FL round, the server broadcasts $\mathbf{W}_{t}$ to all the clients. Then, each client $k$ sets $\mathbf{W}_{t}^{k}=\mathbf{W}_{t}$ and computes the $\tau\;(\geq 1)$ local parameter vectors by SGD: [^2] When $T$ is fixed, the larger $\tau$ is, the fewer the communication rounds there are.
$\mathbf{W}_{t+t^{\prime}}^{k}\leftarrow\mathbf{W}_{t+t^{\prime}-1}^{k}-\eta_{t+t^{\prime}-1}\nabla F_{k}^{j_{t+t^{\prime}-1}^{k}}(\mathbf{W}_{t+t^{\prime}-1}^{k}),\;\;t^{\prime}=1,\ldots,\tau,$ (3)
where $\eta_{t+t^{\prime}-1}$ is the learning rate, $\nabla F_{k}^{j}(\mathbf{W}):=\nabla F_{k}(\mathbf{W};\xi_{k,j})$ is the gradient computed at a single sample of index $j$, and $j_{t}^{k}$ is the sample index chosen uniformly from the local data $\mathcal{D}^{(k)}$ of client $k$ at time tt. [^3] In this work, our focus is on analyzing the computation of a single stochastic gradient at each client during every time instance. The FL convergence rates can potentially be enhanced by incorporating mini-batching technique [stich2018local], leaving the detailed analysis for future work. Finally, by assuming that all clients participate in each FL round for simplicity, the server aggregates the $K$ local model updates $\{\mathbf{X}_{t+\tau}^{k}=\mathbf{W}_{t+\tau}^{k}-\mathbf{W}_{t}\}_{k\in\mathcal{K}}$ and computes the new global parameter vector:
@@
In [shlezinger2020uveqfed], Shlezinger et al. introduced a framework for Quantized Federated Learning. Given the limited bandwidth of uplink channel in FL framework, each $k$-th client is required to communicate a quantized version [gray1998quantization] of the model update $\mathbf{X}_{t+\tau}^{k}$ using a finite number of bits to the server. Specifically, the $k$-th model update $\mathbf{X}_{t+\tau}^{k}$ is encoded into a binary codeword $u_{t+\tau}^{k}$ of length $R_{k}$ bits through an encoding function $\mathrm{Enc}_{t+\tau}^{k}:\mathbb{R}^{m}\to\{0,1,\ldots,2^{R_{k}}-1\}:=\mathcal{U}_{k}$. Since the outputs $\mathrm{Enc}_{t+\tau}^{k}(\mathbf{X}_{t+\tau}^{k})=u_{t+\tau}^{k}\in\mathcal{U}_{k}$ of an encoder may have unequal probabilities, entropy coding [huffman1952method, Golomb1966enc, elias1975universal] can be utilized to further reduce redundancy when transmitting through the binary lossless channel. In the FL literature, the uplink channel is typically modeled as a bit-constrained link [shlezinger2020uveqfed], and the transmission is assumed to be lossless. Upon receiving the set of codewords $\{u_{t+\tau}^{k}\}_{k\in\mathcal{K}}$ from clients, the server uses the joint decoding function $\mathrm{Dec}_{t+\tau}:\mathcal{U}_{1}\times\ldots\times\mathcal{U}_{K}\to\mathbf{R}^{m}$ to reconstruct $\hat{\mathbf{X}}_{t+\tau}\in\mathbb{R}^{m}$, an estimate of the weighted average $\sum_{k\in\mathcal{K}}p_{k}\mathbf{X}_{t+\tau}^{k}$.